OpenStudy (elvisiscool69):

what is an equation of the line in slope intercept form? m=4 and the y-intercept is (0,5) a.y=4x-5 b.y=-5x+4 c.y=4x+5 d.y=5x+4

OpenStudy (dayakar):

do u know slope intercept form

OpenStudy (dayakar):

what is the y- intercept in the problem

OpenStudy (crabbyoldgamer):

Slope intercept form is y=mx+b where m is the slope and b is the y intercept.

OpenStudy (rose4825):

In the equation of a straight line (when the equation is written as "y = mx + b"), the slope is the number "m" that is multiplied on the x, and "b" is the y-intercept, where the line crosses the y-axis. When x = 0, the corresponding y-value is the y-intercept. Since the y-intercept in your question is (0, 5), then the b would be 5, because x = 0 and y = 5. You have been given the m (or slope) of the equation, m = 4. Now you will plug both values into the equation: y = mx + b. Do you now know what the answer is?